# Explainable AI for computational pathology identifies model limitations and tissue biomarkers

### v2

@@ Results
TILs are a well-known prognostic biomarker. We evaluated the necessity and sufficiency of TILs for low-risk predictions in BRCA and SKCM. To test sufficiency, we extracted TIL-positive patches from low-risk specimens and placed them in high-risk specimens. For each high-risk slide, we embedded the TILs from each low-risk slide, and we averaged the model predictions across the low-risk slides to compute the average treatment effect of TILs for each high-risk slide. In high-risk BRCA specimens (n=253, three specimens failed cell detection), the addition of TILs from low-risk specimens decreased the risk by 46% (, paired t-test) from 0.37 (std. dev. 0.20) to 0.20 (std. dev. 0.15). In SKCM (n=67), the addition of TILs significantly decreased risk by 59% (, paired t-test) from 0.60 (std. dev. 0.14) to 0.25 (std. dev. 0.08) (Fig. 4c). To evaluate the necessity of TILs, we removed TIL-positive patches from low-risk specimens and measured the change in predictions. If TILs were necessary, then risk predictions would increase upon removal of TILs. In BRCA (n=254, two specimens failed cell detection), the removal of TILs significantly increased risk by 179% (, paired t-test) from 0.002 (std. dev. 0.001) to 0.005 (std. dev. 0.014). In SKCM (n=67), the removal of TILs increased risk by 98% (, paired t-test) from 0.064 (std. dev. 0.045) to 0.126 (std. dev. 0.123) (Fig. 4d). The removal of TILs did increase risk predictions, but the risk predictions did not reach the level of high-risk slides, suggesting that other features in the WSIs were also driving the low-risk predictions. HIPPO facilitated a quantitative evaluation of the role of TILs on prognosis, providing insights beyond those achievable through the attention mechanism of ABMIL.
@@
Lifileucel is a promising immunotherapy for melanoma that involves isolating TILs from a patient’s tumor, replicating the TILs, and infusing them back into the patient111https://www.fda.gov/news-events/press-announcements/fda-approves-first-cellular-therapy-treat-patients-unresectable-or-metastatic-melanoma. In a phase II clinical trial, over 30% of patients responded to the therapy [89]. Identifying the patients that might respond to this therapy has the potential to improve patient outcomes and decrease costs (a single treatment may cost over $] [90]). Therefore, we sought to explore whether we could emulate this with ABMIL and HIPPO. We conducted virtual experiments to measure the effect of autologous TILs on prognosis. We used the prognostic model for cutaneous melanoma described above, and we studied the high-risk specimens in TCGA-SKCM (n=67 WSIs, n=54 patients). Counterfactuals were designed to model the injection of autologous TILs. In each specimen, TIL-positive patches were replicated , , , and (Supplementary Fig. 11a). TIL-positive patches were defined using the same heuristic as above (see Methods). The change in model predictions between original specimens and autologous counterfactuals was recorded to measure the effect of additional TILs on prognosis. Cohen’s d was also calculated to quantify effect sizes. Importantly, we do not claim to demonstrate the efficacy of autologous TIL therapy through HIPPO and TCGA-SKCM. Rather, we aim to show a proof-of-principle that HIPPO may be used for hypothesis generation.
Autologous TILs significantly lowered predicted risk in a dose-dependent manner. Risk decreased by () at dose (, , paired t-test), () at dose (, , paired t-test), () at dose (, , paired t-test), and () at dose (, , paired t-test) (Supplementary Fig. 11b). Increasing the number of TILs by decreased predicted risk scores by over half in of high-risk specimens. Together, we demonstrated a proof-of-principle in which we use HIPPO to identify patients who may benefit from autologous TIL therapy through improved predicted prognosis following the replication of their TILs.
@@
To evaluate HIPPO’s advantages over attention in diagnosing misclassifications, we applied it to IDH mutation classification, a prognostic marker in gliomas [91]. Using ABMIL models trained on the EBRAINS dataset [92] (n=873), we classified wildtype (n=540) and mutant (n=333) IDH in H&E-stained specimens, employing 10-fold Monte Carlo cross-validation to assess variability. The classifiers demonstrated strong performance, with balanced accuracy of (mean std. dev.) (Fig. 5a), though the false negative (FN) rate ranged from 0.03 to 0.26 across folds. We used HIPPO and attention to identify and remove influential patches in FN specimens, aiming to increase the model’s IDH mutation probability.
@@
When removing the top of patches identified by the strategy HIPPO-search-high-effect, predicted IDH mutation probabilities increased significantly from baseline (, independent t-test, Cohen’s ), resulting in a true positive classification for of cases. In comparison, removing the top of high-attention patches only converted FNs to true positives in of cases, with a nonsignificant effect on mutation probability (, independent t-test, Cohen’s ) (Fig. 5b). These results suggest that HIPPO was more effective than attention at identifying critical regions causing FN predictions.
@@
While there was a large degree of overlap between the patches identified by HIPPO and attention (median: , range: to ), the patches exclusive to HIPPO were often key to resolving FN classifications. A representative sample of these patches is shown in Fig. 5c. Highlighting this, in specimen a1982c61-357f-11eb-b540-001a7dda7111, occluding high-HIPPO patches raised the IDH mutation probability from 0.24 to 0.93 (Fig. 5d), while occluding high-attention patches raised it only to 0.30 (Fig. 5e). Although HIPPO and attention overlapped by , the unique patches identified by HIPPO proved crucial in converting this FN prediction to a true positive prediction (Fig. 5e).
@@
HIPPO also demonstrated clear advantages over attention in identifying regions that drive positive predictions. Using HIPPO-search-high-effect, we identified the patches most strongly contributing to positive predictions and compared these to patches prioritized by attention. HIPPO consistently identified more impactful patches, as shown by the steeper and more substantial decreases in model predictions when HIPPO-identified patches were removed compared to those identified by attention (Supplementary Figs. 12a and 12c).
@@
Heatmaps generated from HIPPO outputs revealed a sharper delineation of critical regions than attention maps, which may enhance interpretability. Interestingly, HIPPO sometimes highlighted different regions than those with high attention scores, suggesting that diagnostically relevant areas may not always correspond to high-attention regions (Supplementary Figs. 12b and 12d). To further assess patch quality, we measured how many patches needed to be removed to lower a positive prediction below a probability threshold of 0.4. HIPPO required significantly fewer patches than attention (, independent t-test), indicating that it identified higher-quality, diagnostically relevant patches.
@@ Discussion
In this study, we introduce HIPPO, an explainable AI method designed to enhance the interpretability and trustworthiness of ABMIL models in computational pathology. Our results demonstrate HIPPO’s ability to uncover hidden biases, quantify the impact of specific tissue regions on model predictions, and bridge the gap between computational outputs and clinically relevant insights. These findings may have significant implications for the development, regulation, and clinical application of AI in pathology.
@@
While our study demonstrates the potential of HIPPO, several limitations must be acknowledged. First, the counterfactual scenarios generated by HIPPO, while informative, may not always reflect biologically plausible tissue alterations. Future work should focus on refining these interventions to more closely mimic realistic tissue changes. Second, HIPPO operates at the resolution of patches, which represents a fundamental limit in its ability to capture finer-grained tissue-level details. Additionally, our analysis was limited to a specific set of foundation models and datasets. Broader evaluation across diverse pathology tasks and model architectures is needed to fully characterize the generalizability of our findings. In addition the interpretations offered by HIPPO are inherently bound by the underlying model’s capabilities and potential shortcomings in representing complex biological systems.
Looking ahead, this work opens several promising directions for future research. Integrating HIPPO with multimodal data, such as genomic and clinical information, could deepen its capacity to reveal insights into both model behavior and biological relevance. While HIPPO identifies important patches de novo, interpreting these patches often requires the expertise of trained pathologists to discern the shared features driving model predictions. Alternatively, observational statistics derived from the patches, such as nuclear pleomorphism from nuclear segmentation or tissue type classifications, could help identify key features within HIPPO-identified patches. Incorporating matched spatial transcriptomics data offers another compelling avenue [93], potentially enabling biomarker discovery by linking sub-visual molecular features with morphological characteristics. Such integrations could help overcome the patch-level resolution limit of HIPPO, offering more granular insights into the interplay between molecular features and patient prognosis.
@@ Methods
HIPPO (Histopathology Interventions of Patches for Predictive Outcomes) is an explainable AI toolkit for attention-based multiple instance learning (ABMIL) models in computational pathology. It generates counterfactual examples by manipulating whole slide image patches to explain model behavior. In ABMIL, tissue from a whole slide image (WSI) is divided into small tiles, which are embedded using a pre-trained model. These patch embeddings serve as input to the ABMIL model, which learns to map sets of patch embeddings to specimen-level labels. HIPPO is made possible by two key features of ABMIL: (1) models are invariant to patch order, and (2) models accommodate variable number of patches. Taking advantages of these features of ABMIL, HIPPO creates counterfactual examples by adding or removing patches, allowing for evaluation of hypothetical scenarios and quantification of the effects of tissue region on model predictions.
When annotations are available, they can guide hypothesis-driven experiments. Patches corresponding to specific annotated regions can be removed or added to test their necessity or sufficiency for model predictions. For instance, in the present study, we utilized expert annotations of tumor regions to investigate the necessity and sufficiency of tumor presence in predicting metastasis. By removing tumor patches, we assessed their impact on model predictions to determine their criticality. Conversely, tumor patches were added to other specimens to evaluate their sufficiency in driving positive predictions. Automated methods may also be used to generate annotations. In the present report, for example, regions positive for tumor-infiltrating lymphocytes were defined using an off-the-shelf cell classification network, and these regions were used to test hypotheses in prognostic models.
@@
Additionally, the HIPPO search algorithm offers an alternative for identifying critical patches. This method directly evaluates the effect of each patch on the model’s predictions, accommodating potential nonlinear interactions among patches. The algorithm operates as follows:
@@
Iteratively drop one patch at a time, measure the change in the model’s prediction, and replace the patch.
@@
Identify and permanently remove the patch that results in the lowest prediction (to identify necessary patches) or the smallest change (to identify sufficient patches).
@@
Repeat the process until all patches are exhausted or a stopping criterion is met.
The HIPPO search algorithm provides a systematic framework for identifying patches that are most critical to model predictions. This can be particularly useful when hypothesis-driven experiments reveal discrepancies between expected and actual model behavior, or when the important tissue features are not known beforehand. By identifying the regions the model prioritizes, this approach enables comparisons with domain knowledge, helps uncover unexpected decision-making patterns, and may aid in biomarker discovery. A limitation of the HIPPO search strategy is potentially long runtime, depending on the number of patches in a specimen. To address this, a batched approach can be used, wherein multiple patches are removed at each iteration. By adjusting the batch size, a trade-off can be achieved between computational efficiency and the granularity of the results.
@@
We used attention-based multiple instance learning (ABMIL) to learn specimen-level labels from wholes slide images. For metastasis detection, we evaluated five different patch encoders: UNI [20], REMEDIS [16], CTransPath [17], Phikon [18], and RetCCL [19]. These encoders were used to embed non-overlapping patches, with all encoders utilizing identical patches. We standardized hyperparameters across all ABMIL models, adapting from Chen et al. [20]. The architecture comprised a first hidden layer of 512 units and a second of 384 units, incorporating gated attention. During training, we applied a dropout rate of 0.25. The output layer performed binary classification, distinguishing between the presence and absence of metastasis. Models were trained using cross-entropy loss and the Adam optimizer with a learning rate of , following a cosine learning rate scheduler. We used a batch size of 1 without gradient accumulation. Training continued for a maximum of 20 epochs, with the best model selected based on the highest ROC AUC on the validation set. To assess initialization variability, we trained five separate models with different random seeds for each patch encoder. For subsequent experiments, we selected the initialization yielding the highest balanced accuracy on the CAMELYON16 test set for each encoder. We visualized attention heatmaps using QuPath [94]. All models were implemented in PyTorch and trained on NVIDIA RTX 2080 Ti GPUs.
The setup for IDH mutation classification was similar to that of metastasis classification. However, CONCH [21] embeddings were used, and 10-fold Monte Carlo cross validation was employed to assess variability across models. All models were implemented in PyTorch and trained on NVIDIA RTX 2080 Ti GPUs.
For prognostic models, we used the ABMIL models defined in [50]. The model was composed of a linear layer with 512 units, dropout with a rate of 0.25, and a second linear layer of 256 units. Gated attention was used. The model had four outputs, representing hazards at four points in time. Risk scores were calculated as in ref. [50] and were in range , where 0 indicates lowest probability of survival. Models were all implemented in PyTorch, and training was performed on NVIDIA RTX 2080 Ti GPUs.
@@
We used the CAMELYON16 dataset [30] to study breast cancer metastasis. This dataset consists of 399 images and has fine-grained tumor annotations made by expert pathologists. The training set was split into 90% training and 10% validation, stratified by the label of the specimen (i.e., normal or tumor). Training set consisted of 143 negative and 100 positive WSIs (52 macrometastases and 48 micrometastases). The validation set consisted of 16 negative and 11 positive WSIS (6 macrometastases and 5 micrometastases). We used the pre-defined test set, which consisted of 80 negative and 49 positive WSIs (22 macrometastases and 27 micrometastases). In the entire dataset, there were 160 metastasis-positive specimens. There was an average tumor area of (std. dev. ; minimum ; and maximum ). All 399 slides had pixel spacings between nd (MPP). The WSIs had patches (mean standard deviation), where each patch was .
@@
Prognostic models were trained and evaluated using the invasive breast carcinoma (BRCA) and cutaneous melanoma (SKCM) studies from The Cancer Genome Atlas. In TCGA BRCA, 1,022 WSIs from 956 patients were used (130 death events), and in TCGA SKCM, 268 slides from 230 patients were used (89 death events). Overall survival time and censoring was used and retrieved from the code repository222https://github.com/mahmoodlab/PORPOISE of ref. [50]. The training and validation splits for cross validation were accessed from the same code repository. The WSIs in TCGA BRCA had patches (mean standard deviation). The WSIs in TCGA SKCM had patches.
@@
The EBRAINS dataset [92], comprising 795 patients and 873 specimens with known IDH mutation status, was used to develop IDH mutation classification models. The dataset included 540 IDH wildtype and 333 IDH mutant specimens, spanning 508 glioblastomas, 189 astrocytomas, and 176 oligodendrogliomas. All WSIs were in NDPI format with physical spacing of . Specimens were partitioned for 10-fold Monte Carlo cross-validation using the CLAM toolkit [62]. The WSIs had patches (mean standard deviation), where each patch was .
@@
Whole slide images were read using OpenSlide [95], and a modified version of the CLAM toolkit [62] was used to segment tissue and calculate patch coordinates. Regions of tissue are identified to not spend computational resources on glass regions of the WSI. The image is converted to HSV color model (hue, saturation, value/brightness). The saturation channel is smoothed and thresholded to create a binary tissue image. Non-overlapping patches coordinates of were calculated within the tissue regions. The CLAM toolkit [62] was modified to create patches at uniform physical sizes. The size of a patch in pixels can vary based on the spacing (, MPP), and the patch size in base pixels is calculated using,
The patches from CAMELYON16 were then embedded using five pre-trained models (embedding dimensions in parentheses): UNI (1024) [20], REMEDIS (4096) [16], Phikon (768) [18], CTransPath (768) [17], and RetCCL (2048) [19]. The TCGA-BRCA and TCGA-SKCM specimens for prognostic models were embedded using the UNI model, and the glioma specimens from EBRAINS were embedded using the CONCH model. When embedding a slide, patches were read directly from the WSI file. A batch of 64 patches was read from the WSI, and then that batch was processed by the model to compute embeddings. The embeddings of all patches in a WSI were concatenated into one array and saved to disk for reuse. These embeddings served as inputs to all ABMIL models in the present report.

### v3

@@ Results
TILs are a well-established prognostic biomarker. We evaluated their necessity and sufficiency for low-risk predictions in BRCA and SKCM using HIPPO. To test sufficiency, we extracted TIL-positive patches from low-risk specimens and placed them in high-risk specimens (Supplementary Fig. LABEL:fig:hippo-schematic-insert-deleteb). For each high-risk slide, we averaged predictions across TILs sourced from low-risk slides to compute the average treatment effect. In high-risk BRCA specimens (n=253, three specimens failed cell detection), adding TILs from low-risk specimens decreased the risk by 46% (, paired two-sided t-test) from 0.37 (std. dev. 0.20) to 0.20 (std. dev. 0.15). In SKCM (n=67), TIL addition significantly decreased risk by 59% (, paired two-sided t-test) from 0.60 (std. dev. 0.14) to 0.25 (std. dev. 0.08) (Fig. 4c).
To evaluate the necessity of TILs, we removed TIL-positive patches from low-risk specimens and measured the change in predictions (Supplementary Fig. LABEL:fig:hippo-schematic-insert-deletea). If TILs are necessary for low-risk predictions, risk should increase upon removal. In BRCA (n=254, two excluded), removaling TILs increased risk by 179% (, paired two-sided t-test) from 0.002 (std. dev. 0.001) to 0.005 (std. dev. 0.014). In SKCM (n=67), risk increased by 98% (, paired two-sided t-test) from 0.064 (std. dev. 0.045) to 0.126 (std. dev. 0.123) (Fig. 4d). Although removing TILs increased risk, the resulting risk predictions did not reach the level of high-risk slides, suggesting that other features in the WSIs were also driving the low-risk predictions. Together, these analyses show that HIPPO facilitates a quantitative evaluation of the role of prognostic features, such as TILs, providing insights not achievable through attention alone.
@@
Lifileucel is a promising immunotherapy for melanoma that involves isolating TILs from a patient’s tumor, expanding them ex vivo, and reinfusing them into the patient111https://www.fda.gov/news-events/press-announcements/fda-approves-first-cellular-therapy-treat-patients-unresectable-or-metastatic-melanoma. In a phase II clinical trial, more than 30% of patients responded to the treatment [sarnaik2021lifileucel]. Identifying which patients are most likely to benefit from this therapy could meaningfully improve patient outcomes and reduce costs, given that a single treatment may cost over $] [healey2024lifileucelnews]). Motivated by this, we investigated whether HIPPO could be used to generate hypotheses about patient-specific responses using virtual experiments.
We used the prognostic ABMIL model for cutaneous melanoma described above, and we studied the high-risk specimens from TCGA-SKCM (n=67 WSIs, n=54 patients). Counterfactuals were constructed to emulate autologous TIL infusion: in each specimen, TIL-positive patches were replicated , , , and (Supplementary Fig. LABEL:fig:supp_surv_autologous_tilsa). TIL-positive patches were defined using the heuristic described in Methods. Changes in model-predicted risk between original specimen and each counterfactual was recorded, and Cohen’s d was calculated to quantify effect sizes. Importantly, we emphasize that these experiments are not intended to evaluate the true efficacy of autologous TIL therapy, but rather to illustrate how HIPPO can be used for hypothesis generation.
Autologous TIL augmentation reduced predicted risk in a dose-dependent manner. Risk decreased by () at dose (, , paired two-sided t-test), () at dose (, , paired two-sided t-test), () at dose (, , paired two-sided t-test), and () at dose (, , paired two-sided t-test) (Supplementary Fig. LABEL:fig:supp_surv_autologous_tilsb). Increasing the number of TILs by decreased predicted risk scores by over half in of high-risk specimens. Together, these results demonstrate a proof-of-principle that HIPPO can be used to generate hypotheses about which patients may derive prognostic benefit from autologous TIL therapy, by simulating the effect of TIL replication on patient-specific risk.
@@
To evaluate HIPPO’s advantages over attention in diagnosing misclassifications, we applied it to IDH mutation classification, a prognostic marker in gliomas [han2020idh]. Using ABMIL models trained on the EBRAINS dataset [roetzer2022digital] (n=873), we classified wildtype (n=540) and mutant (n=333) IDH in H&E-stained specimens, employing 10-fold Monte Carlo cross-validation to assess variability. The classifiers demonstrated strong performance, achieving a balanced accuracy of (mean std. dev.) (Fig. 5a), although false negative (FN) rates varied widely across folds (0.03-0.26). We used HIPPO and attention to identify and remove influential patches in FN specimens, with the goal of increasing the model’s IDH mutation probability.
@@
Removing the top of patches identified by HIPPO-search-negative-effect significantly increased predicted IDH mutation probabilities from baseline (, independent two-sided t-test, Cohen’s ), converting of FNs into true positives. In contrast, removing the top of high-attention patches converted only of FNs, producing a nonsignificant change in mutation probability (, independent two-sided t-test, Cohen’s ) (Fig. 5b). These results indicate that HIPPO is more effective than attention at identifying critical regions responsible for FN predictions.
@@
While there was substantial overlap between the patches identified by HIPPO and attention (median: , range: to ), the patches uniquely identified by HIPPO were often critical for correcting FNs (Fig. 5c). For instance, in specimen a1982c61-357f-11eb-b540-001a7dda7111, occluding high-HIPPO patches increased the predicted IDH mutation probability from 0.24 to 0.93 (Fig. 5d), whereas occluding high-attention patches increased it only to 0.30 (Fig. 5e). Despite an overlap between the two patch sets, the HIPPO-exclusive patches were essential for converting this FN to a true positive (Fig. 5e).
@@
HIPPO also demonstrated clear advantages over attention in identifying regions that drive positive predictions. For specimens predicted as IDH-mutant, applying HIPPO-search-positive-effect consistently revealed patches with stronger contributions to the positive prediction. Removing these HIPPO-identified patches produced more pronounced decreases in mutation probability than removing high-attention patches (Supplementary Figs. LABEL:fig:supp_ebrains_posa and LABEL:fig:supp_ebrains_posc).
@@
Heatmaps derived from HIPPO outputs exhibited sharper delineation of critical regions compared to attention maps, potentially improving interpretability. Notably, HIPPO sometimes highlighted regions distinct from those with high attention scores, implying that diagnostically relevant areas may not always align with high-attention regions (Supplementary Figs. LABEL:fig:supp_ebrains_posb and LABEL:fig:supp_ebrains_posd). To quantify patch quality, we measured how many patches needed to be removed to reduce positive predictions below a probability threshold of 0.4. HIPPO required significantly fewer patches than attention (, independent two-sided t-test), indicating that HIPPO identified patches with a stronger causal influence on model predictions.
HIPPO is generally applicable to any permutation-invariant MIL models beyond ABMIL
While HIPPO is a versatile toolkit applicable to any permutation-invariant MIL models, our primary experiments concentrated on ABMIL architectures due to their status as the de facto standard in computational pathology. However, since alternative MIL architectures, particularly those based on transformer self-attention, are emerging, we also experimentally demonstrate that HIPPO generalizes beyond ABMIL.
To this end, we extended the IDH mutation analysis to a vision transformer (ViT)-based MIL model [dosovitskiy_image_2021, wagner2023transformer, el2025stamp]. The original ViT model [dosovitskiy_image_2021] was adapted to the MIL setting by using patch embeddings as tokens and removing positional embeddings to ensure permutation invariance [wagner2023transformer, el2025stamp]. Following the evaluation setup in Figure 5b, we assess if HIPPO provides more accurate identification of influential patches than ViT’s intrinsic attention mechanism (Supplementary Fig. LABEL:fig:supp_idh_prediction_with_vit). Specifically, we compared HIPPO-search-negative with two scores derived from ViT’s attention mechanism, namely attention last and attention rollout [abnar2020quantifying] (see Methods).
Removing the top of patches identified by HIPPO-search-negative-effect significantly increased predicted IDH mutation probabilities from baseline (, independent two-sided t-test, Cohen’s ), resulting in a true positive classification for of cases. In comparison, removing the top of ViT-based attention patches converted only of FNs to true positives when using attention last and when using ViT attention rollout, with both producing nonsignificant effects on mutation probability (, independent two-sided t-test, Cohen’s and , respectively). These findings show that HIPPO generalizes effectively to other permutation-invariant MIL architectures beyond ABMIL.
@@ Discussion
In this study, we introduce HIPPO, an explainable AI method designed to enhance the interpretability and trustworthiness of weakly supervised multiple instance learning models in computational pathology. Our results demonstrate HIPPO’s ability to uncover hidden biases, quantify the impact of specific tissue regions on model predictions, and bridge the gap between computational outputs and clinically relevant insights. HIPPO offers a generalizable framework for interpreting, auditing, and stress-testing MIL models, enabling rigorous evaluation and greater confidence in pathology AI systems.
@@
While our study demonstrates the potential of HIPPO, several limitations must be acknowledged. First, the counterfactual scenarios generated by HIPPO, while informative, may not always reflect biologically plausible tissue alterations. Future work should focus on refining these interventions to more closely mimic realistic tissue changes. Second, HIPPO operates at the resolution of patches, which limits its ability to capture finer-grained tissue-level details. Third, although we demonstrate HIPPO’s applicability beyond ABMIL using a permutation-invariant ViT model, HIPPO currently depends on architectures that operate on patch-level embeddings and do not use spatial information. Extending HIPPO to end-to-end or non-MIL models, including foundation models with spatial reasoning or slide-level supervision, remains an important direction for future research.
Additionally, our analysis was limited to a specific set of foundation models and datasets. Broader evaluation across diverse pathology tasks and model architectures is needed to fully characterize HIPPO’s capability in real-world settings as well as the generalizability of our findings. Importantly, HIPPO’s interpretive power is ultimately constrained by the representational capacity of the underlying AI model. Because HIPPO generates counterfactual predictions using the model’s learned associations, any limitations in the model’s ability to capture complex biological relationships will naturally be reflected in its explanations. As a result, HIPPO is best viewed as a rigorous framework for understanding model behavior and generating testable hypotheses, rather than a standalone tool for definitive biomarker discovery in settings where the underlying model’s predictive power is not yet robustly established.
Looking ahead, this work opens several promising directions for future research. Integrating HIPPO with multimodal data, such as genomic and clinical information, could deepen its capacity to reveal insights into both model behavior and biological relevance. While HIPPO identifies important patches de novo, interpreting these patches often requires the expertise of trained pathologists to discern the shared features driving model predictions. Alternatively, observational statistics derived from the patches, such as nuclear pleomorphism from nuclear segmentation or tissue type classifications, could help identify key features within HIPPO-identified patches. Incorporating matched spatial transcriptomics data offers another compelling avenue [jaume2024hest], potentially enabling biomarker discovery by linking sub-visual molecular features with morphological characteristics. Such integrations could help overcome the patch-level resolution limit of HIPPO, offering more granular insights into the interplay between molecular features and patient prognosis.
@@ Methods
HIPPO (Histopathology Interventions of Patches for Predictive Outcomes) is an explainable AI toolkit for multiple instance learning (MIL) models in computational pathology, including attention-based multiple instance learning (ABMIL)[ilse2018abmil] and VisionTransformer models [el2025stamp]. In computational histopathology, tissue from a whole slide image (WSI) is divided into small patches, which are embedded using a pre-trained model. These patch embeddings serve as input to the MIL model, which learns to map sets of patch embeddings to specimen-level labels. HIPPO is made possible by two key features of MIL: (1) models are invariant to patch order, and (2) models accommodate a variable number of patches. Taking advantage of these features of MIL, HIPPO creates counterfactual examples by adding or removing patches, allowing for the evaluation of hypothetical scenarios and the quantification of the effects of tissue region on model predictions. In other words, HIPPO simulates counterfactual WSIs by subsampling patches.
Within the HIPPO framework, users can flexibly select which patches to include or exclude, depending on their research question and settings. When annotations are available, one can consider the HIPPO-knowledge approach , enabling hypothesis-driven experiments. Patches corresponding to specific annotated regions can be removed or added to test their necessity or sufficiency for model predictions. For instance, in the present study, we utilized expert annotations of tumor regions to investigate the necessity and sufficiency of tumor presence in predicting metastasis. By removing tumor patches, we assessed their impact on model predictions to determine their criticality. Conversely, tumor patches were added to other specimens to evaluate their sufficiency in driving positive predictions. Automated methods may also be used to generate annotations. In the present report, for example, regions positive for tumor-infiltrating lymphocytes were defined using an off-the-shelf cell classification network, and these regions were used to test hypotheses in prognostic models.
@@
Additionally, the HIPPO search algorithm offers an alternative for identifying critical patches de novo. This method directly evaluates the effect of each patch on the model’s predictions, accommodating potential nonlinear interactions among patches. The algorithm operates as follows:
@@
Iteratively, for each remaining patch in the WSI, we temporarily drop one patch and measure the resulting change in the model’s prediction.
@@
Identify and permanently remove the patch whose impact of removal is strongest. To do so, we first compute the effect size of each patch, which is defined as the original model output minus the model output after the patch is removed. For a specimen predicted as positive, we identify the patch with the most positive effect size (argmax). For a specimen predicted as negative, we identify the patch with the most negative effect size (argmin).
@@
Repeat the process until all patches are exhausted or a stopping criterion is met (e.g., a certain number or percentage of patches removed) .
The HIPPO search algorithm provides a systematic framework for identifying patches that are most critical to model predictions. This can be particularly useful when hypothesis-driven experiments reveal discrepancies between expected and actual model behavior, or when the important tissue features are not known beforehand. By identifying the regions the model prioritizes from scratch, this approach enables comparisons with domain knowledge, helps uncover unexpected decision-making patterns, and may aid in biomarker discovery. A limitation of the HIPPO search strategy is potentially long runtime, depending on the number of patches in a specimen. To address this, a batched approach can be used, wherein multiple patches are removed at each iteration. By adjusting the batch size, a trade-off can be achieved between computational efficiency and the granularity of the results. Detailed algorithmic descriptions with pseudo-code for all HIPPO usages (HIPPO-knowledge, HIPPO-attention, HIPPO-search-positive-effect, and HIPPO-search-negative-effect) are provided in the LABEL:sec:supplementary_methods.
@@
We used attention-based multiple instance learning (ABMIL) to learn specimen-level labels from wholes slide images. For metastasis detection, we evaluated five different patch encoders: UNI [chen2024uni], REMEDIS [azizi2022remedis], CTransPath [wang2022ctranspath], Phikon [filiot2023phikon], RetCCL [wang2023retccl], and Virchow2 [vorontsov2024virchow, zimmermann2024virchow2]. These encoders were used to embed non-overlapping patches, with all encoders utilizing identical patches. We standardized hyperparameters across all ABMIL models, adapting from Chen et al. [chen2024uni]. The architecture comprised a first hidden layer of 512 units and a second of 384 units, incorporating gated attention. During training, we applied a dropout rate of 0.25. The output layer performed binary classification, distinguishing between the presence and absence of metastasis. Models were trained using cross-entropy loss and the Adam optimizer with a learning rate of , following a cosine learning rate scheduler. We used a batch size of 1 without gradient accumulation. Training continued for a maximum of 20 epochs, with the best model selected based on the highest ROC AUC on the validation set. To assess initialization variability, we trained five separate models with different random seeds for each patch encoder. For subsequent experiments, we selected the initialization yielding the highest balanced accuracy on the CAMELYON16 test set for each encoder. We visualized attention heatmaps using QuPath [bankhead2017qupath]. All models were implemented in PyTorch and trained on NVIDIA RTX 2080 Ti GPUs.
The setup for IDH mutation classification was similar to that of metastasis classification. However, CONCH [lu2024visualconch] embeddings were used, and 10-fold Monte Carlo cross validation was employed to assess variability across models. All models were implemented in PyTorch and trained on NVIDIA RTX 2080 Ti GPUs.
For prognostic models, we used the ABMIL models defined in [chen2022porpoise]. The model was composed of a linear layer with 512 units, dropout with a rate of 0.25, and a second linear layer of 256 units. Gated attention was used. The model had four outputs, representing hazards at four points in time. Risk scores were calculated as in ref. [chen2022porpoise] and were in range , where 0 indicates lowest probability of survival. Models were all implemented in PyTorch, and training was performed on NVIDIA RTX 2080 Ti GPUs.
@@
We used the CAMELYON16 dataset [bejnordi2017camelyon16] to study breast cancer metastasis. This dataset consists of 399 images and has fine-grained tumor annotations made by expert pathologists. These expert-annotated tumor segmentations provide ground-truth labels that enable systematic validation of HIPPO’s causal attributions against known biological features. Most of our HIPPO experiments rely on these expert-annotated ground-truth tumor regions to validate the necessity and sufficiency of tissue regions for model predictions. The training set was split into 90% training and 10% validation, stratified by the label of the specimen (i.e., normal or tumor). Training set consisted of 143 negative and 100 positive WSIs (52 macrometastases and 48 micrometastases). The validation set consisted of 16 negative and 11 positive WSIS (6 macrometastases and 5 micrometastases). We used the pre-defined test set, which consisted of 80 negative and 49 positive WSIs (22 macrometastases and 27 micrometastases). In the entire dataset, there were 160 metastasis-positive specimens. There was an average tumor area of (std. dev. ; minimum ; and maximum ). All 399 slides had pixel spacings between nd (MPP). The WSIs had patches (mean standard deviation), where each patch was .
@@
Prognostic models were trained and evaluated using the invasive breast carcinoma (BRCA) and cutaneous melanoma (SKCM) studies from The Cancer Genome Atlas. In TCGA BRCA, 1,022 WSIs from 956 patients were used (130 death events), and in TCGA SKCM, 268 slides from 230 patients were used (89 death events). Overall survival time and censoring was used and retrieved from the code repository222https://github.com/mahmoodlab/PORPOISE of ref. [chen2022porpoise]. The training and validation splits for cross validation were accessed from the same code repository. The WSIs in TCGA BRCA had patches (mean standard deviation). The WSIs in TCGA SKCM had patches.
@@
The EBRAINS dataset [roetzer2022digital], comprising 795 patients and 873 specimens with known IDH mutation status, was used to develop IDH mutation classification models. The dataset included 540 IDH wildtype and 333 IDH mutant specimens, spanning 508 glioblastomas, 189 astrocytomas, and 176 oligodendrogliomas. All WSIs were in NDPI format with physical spacing of . Specimens were partitioned for 10-fold Monte Carlo cross-validation using the CLAM toolkit [lu2021clam]. The WSIs had patches (mean standard deviation), where each patch was .
@@
Whole slide images were read using OpenSlide [goode2013openslide], and a modified version of the CLAM toolkit [lu2021clam] was used to segment tissue and calculate patch coordinates. Regions of tissue are identified to not spend computational resources on glass regions of the WSI. The image is converted to HSV color model (hue, saturation, value/brightness). The saturation channel is smoothed and thresholded to create a binary tissue image. Non-overlapping patches coordinates of were calculated within the tissue regions. The CLAM toolkit [lu2021clam] was modified to create patches at uniform physical sizes. The size of a patch in pixels can vary based on the spacing (, MPP), and the patch size in base pixels is calculated using,
The patches from CAMELYON16 were then embedded using six pre-trained models (embedding dimensions in parentheses): UNI (1024) [chen2024uni], REMEDIS (4096) [azizi2022remedis], Phikon (768) [filiot2023phikon], CTransPath (768) [wang2022ctranspath], RetCCL (2048) [wang2023retccl], and and Virchow2 (1280) [vorontsov2024virchow, zimmermann2024virchow2]. The TCGA-BRCA and TCGA-SKCM specimens for prognostic models were embedded using the UNI model, and the glioma specimens from EBRAINS were embedded using the CONCH model. When embedding a slide, patches were read directly from the WSI file. A batch of 64 patches was read from the WSI, and then that batch was processed by the model to compute embeddings. The embeddings of all patches in a WSI were concatenated into one array and saved to disk for reuse. These embeddings served as inputs to all ABMIL models in the present report.